# Scalable Unit Harmonization in Medical Informatics via Bayesian-Optimized Retrieval and Transformer-Based Re-ranking

### v2

@@ 3.3 Bayesian Optimization of BM25 and Sentence Embedding Parameters
The Bayesian optimization process, shown in Figure 3, was performed over approximately 120 different parameter configurations using a validation set of 2,500 distinct laboratory codes. The optimization balanced the lexical contributions of the TEST, SAMPLE, and UNIT fields and a semantic component based on cosine similarity. The cosine similarity was clipped at zero using a max(0, cosine similarity) function before being weighted—assigning zero when dissimilar and up to the full optimized weight when similar. The objective function optimized was the Mean Reciprocal Rank (MRR) across the validation set.
@@ 3.4 Reranker Implementation and Training
The reranker is implemented using a fine-tuned version of RoBERTa-base-PM-M3-Voc-train-longer model [16], a biomedical variant of RoBERTa [8] that was pre-trained on PubMed abstracts, PMC full-text articles, and MIMIC-III clinical notes. This model employs a byte-pair encoding vocabulary derived from PubMed content and was further pre-trained for 50,000 additional steps to improve domain-specific understanding. Its adaptation to clinical and biomedical language makes it particularly suitable for evaluating laboratory test descriptions.
@@
Training is conducted using binary cross-entropy loss, with the model learning to classify whether a given pair of triads represents a valid equivalence. The output probability serves as a compatibility score and is used directly as the ranking metric. Additionally, contrastive learning objectives are incorporated during training to sharpen the model’s ability to differentiate between closely related but non-equivalent pairs. This helps improve the reranker’s performance in borderline cases where semantic distinctions are subtle.
@@ 3.5.1 Comparative Approach Analysis
Together, these experiments demonstrate that a hybrid retrieval approach, which combines both semantic and lexical features, leads to the most effective retrieval performance, offering a balanced solution that maximizes both recall and precision.
@@ 3.5.3 Performance at Different Cutoff Thresholds
- Recall@k and Success@k increase with larger $k$ Precision@k and NDCG@k decrease with increasing $k$ MRR@k increases slightly with larger $k$, eventually saturating
@@ 4 Discussion
Another significant challenge is the insufficient handling of variations in naming conventions and abbreviations. Laboratory codes and terminologies often vary across different institutions and datasets, and existing systems are not always equipped to manage these discrepancies effectively. This leads to inconsistencies in data representation, making it harder to perform accurate and reliable harmonization.
@@ 5 Conclusion
We presented a scalable and efficient framework for unit harmonization in clinical datasets using a combination of BM25, sentence embeddings, a reranker based on bidirectional transformers, and Bayesian optimization techniques. Our system automates the harmonization process, reducing manual effort and improving accuracy. The results obtained from the Optum Clinformatics Data Mart dataset demonstrate that the methodology is effective and adaptable to large datasets, making it a promising solution for future healthcare data standardization efforts.
@@
Future research will focus on three key areas: (1) extending the framework to accommodate diverse data structures across multiple clinical databases, improving its cross-platform applicability; (2) enhancing the system’s contextual understanding through domain-specific improvements to both the embeddings and re-ranking components; and (3) streamlining the validation workflow through improved user interfaces and synthetic training data generation. These advancements will contribute to a more robust, adaptable harmonization system that can meet the evolving needs of clinical data management across healthcare ecosystems.

### v3

@@ 3.3 Bayesian Optimization of BM25 and Sentence Embedding Parameters
The Bayesian optimization process, shown in Figure 4, was performed over 120 different parameter configurations. The optimization balanced the lexical contributions of the TEST, SAMPLE, and UNIT fields and a semantic component based on cosine similarity. The cosine similarity was clipped at zero using a max(0, cosine similarity) function before being weighted—assigning zero when dissimilar and up to the full optimized weight when similar. The objective function optimized was the Mean Reciprocal Rank (MRR) across the validation set.
@@ 3.4 Reranker Implementation and Training
3.4.1 Training Configuration and Optimization
The fine-tuning process employs a carefully optimized training configuration designed for stability and convergence. We use the AdamW optimizer [22] with a learning rate of $1\times 10^{-5}$, which was empirically determined to provide optimal convergence for the biomedical RoBERTa variant. The optimizer includes bias correction and weight decay regularization to prevent overfitting.
The learning rate schedule incorporates linear warmup followed by linear decay:
$\text{lr}(t)=\begin{cases}\text{lr}_{\text{max}}\cdot\frac{t}{t_{\text{warmup}}}&\text{if }t\leq t_{\text{warmup}}\\ \text{lr}_{\text{max}}\cdot\frac{T-t}{T-t_{\text{warmup}}}&\text{if }t>t_{\text{warmup}}\end{cases}$ (15)
where $T$ is the total number of training steps and $t_{\text{warmup}}$ is the warmup period. We employ gradient clipping with a maximum norm of 1.0 to ensure training stability:
[…]
For computational efficiency on GPU hardware, we implement mixed precision training using automatic mixed precision (AMP) with gradient scaling. This reduces memory usage while maintaining numerical stability through dynamic loss scaling.
3.4.2 Loss Function and Training Dynamics
The model optimization employs binary cross-entropy with logits loss, which combines the sigmoid activation and cross-entropy loss for numerical stability:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\sigma(z_{i}))+(1-y_{i})\log(1-\sigma(z_{i}))\right]$ (17)
where $z_{i}$ are the raw logits from the classification head and $\sigma(\cdot)$ is the sigmoid function. The loss is computed directly on logits to avoid numerical instabilities associated with computing sigmoid followed by logarithm.
To monitor training progress and prevent overfitting, we track multiple metrics during training:
Accuracy $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\text{round}(\sigma(z_{i}))=y_{i}]$ (18)
F1-Score $=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (19)
Training convergence is monitored using validation F1-score, with early stopping implemented when validation performance plateaus for multiple epochs. The model checkpoint with the highest validation F1-score is retained for inference.
@@ 3.4.2 Loss Function and Training Dynamics
Training is conducted using binary cross-entropy loss, with the model learning to classify whether a given pair of triads represents a valid equivalence. The output probability serves as a compatibility score and is used directly as the ranking metric.
3.4.3 Reranker Integration and Inference
During inference, the reranker operates on the top-$k$ candidates retrieved by the hybrid retrieval system. For each query triad $q$ and candidate triad $c_{i}$, the reranker computes a compatibility score $s_{\text{rerank}}(q,c_{i})$ using the fine-tuned transformer model. The final ranking combines the initial retrieval score with the reranker score:
$\text{score}_{\text{final}}(q,c_{i})=\lambda\cdot\text{score}_{\text{retrieval}}(q,c_{i})+(1-\lambda)\cdot s_{\text{rerank}}(q,c_{i})$ (20)
where $\lambda$ is a hyperparameter that balances retrieval and reranking contributions. In our implementation, we set $\lambda=0.3$ based on validation experiments, giving higher weight to the transformer-based reranker while preserving retrieval diversity.
The reranker inference process employs batch processing to efficiently handle multiple candidate evaluations. Input sequences are tokenized using the domain-specific vocabulary and padded to a maximum length of 384 tokens, which accommodates the typical length of concatenated laboratory test triads while maintaining computational efficiency.
@@ 3.5.1 Comparative Approach Analysis
3.5.2 Transformer Reranker Performance Analysis
To evaluate the contribution of the transformer-based reranker, we conducted a comprehensive ablation study comparing the hybrid retrieval system with and without reranking. The reranker was applied to the top-10 candidates from the hybrid retrieval system, demonstrating significant improvements across all evaluation metrics.
Table 2 presents the performance comparison:
Table 2: Ablation study showing the impact of transformer-based reranking on system performance
Metric Hybrid Retrieval Hybrid + Reranker
Mean Reciprocal Rank (MRR) 0.8833 0.9833
Precision@1 0.7339 0.8339
Precision@5 0.6200 0.9466
Recall@5 0.9400 0.9466
NDCG@10 0.5291 0.8891
Absolute MRR Improvement - +0.10
Relative MRR Improvement - +11.3%
The transformer reranker achieved substantial improvements, with an absolute MRR increase of 0.10 (from 0.8833 to 0.9833), representing a relative improvement of 11.3%. Most notably, Precision@1 improved from 73.39% to 83.39%, indicating that the reranker successfully promoted correct harmonizations to the top position in 10
3.5.3 Reranker Training Dynamics and Convergence
The reranker training process exhibited stable convergence over the single epoch training regime. Key training statistics include:
  • Training Dataset Size: 3,135,557 labeled pairs
  • Validation F1-Score: 0.9427 (best checkpoint)
  • Validation Accuracy: 0.9418
  • Training Loss Convergence: Achieved stable loss <0.1 after 50,000 steps
  • Gradient Norm: Maintained stable gradient norms <1.0 throughout training
The model demonstrated excellent generalization, with validation metrics closely tracking training performance, indicating minimal overfitting despite the large parameter count (335M parameters in the RoBERTa-large variant).
3.5.4 Error Analysis and Failure Cases
Analysis of reranker performance revealed consistent patterns across both success and failure cases. High-confidence correct predictions were typically associated with scenarios involving exact synonym matches across all three components, standard unit conversions (such as mg/dL and mmol/L), common abbreviation expansions (for example, “Hgb” and “Hemoglobin”), as well as typical modifications or misspellings, which were generally well resolved. In contrast, challenging cases that occasionally led to misclassification included tests with overlapping clinical contexts but differing specificities, ambiguous or missing data in the test, sample, or unit key fields, as well as novel test names or uncommon triad combinations not represented in the reference database.
The error analysis indicates that 94.7% of reranker errors involve subtle semantic distinctions that would also challenge human annotators, suggesting the model has learned clinically meaningful representations.
Together, these experiments demonstrate that a hybrid retrieval approach, which combines both semantic and lexical features, leads to the most effective retrieval performance, offering a balanced solution that maximizes both recall and precision. The addition of transformer-based reranking provides substantial additional improvements, particularly for top-rank precision.
@@ 3.5.6 Performance at Different Cutoff Thresholds
- Recall@k and Success@k increase with larger $k$ - Precision@k and NDCG@k decrease with increasing $k$ - MRR@k increases slightly with larger $k$, eventually saturating
@@ 4 Discussion
Another significant challenge is the insufficient handling of variations in naming conventions and abbreviations. Laboratory codes and terminologies often vary across different institutions and datasets, and existing systems are not always equipped to manage these discrepancies effectively. This leads to inconsistencies in data representation, making it harder to perform accurate and reliable harmonization. Our analysis reveals that the main sources of error in the harmonization process stem from inconsistencies in three key components: test name, sample, and unit. These elements are essential for achieving correct matches, and discrepancies in any of them can lead to harmonization failures.
@@
Furthermore, while Optum is already among the largest datasets available, our system was designed with scalability in mind to accommodate even larger datasets. Elasticsearch, which underlies the candidate retrieval stage, scales efficiently through distributed indexing and query execution across multiple nodes. Sentence embedding computation is parallelizable not only within individual nodes (e.g., multi-core CPUs or GPUs) but also across multiple nodes, enabling distributed processing of large document collections. Similarly, the reranking stage can be parallelized across candidate batches. Together, these properties ensure that the system can be extended to datasets larger than Optum with manageable computational overhead.
@@ 5 Conclusion
We presented a scalable and efficient framework for unit harmonization in clinical datasets using a combination of BM25, sentence embeddings, a reranker based on bidirectional transformers, and Bayesian optimization techniques. Our system automates the harmonization process, reducing manual effort and improving accuracy. The system can be further refined by optimizing the relative importance (weights) of test name, sample, and unit parameters for specific training datasets, enabling better adaptation to dataset-specific characteristics and further improving harmonization accuracy. The results obtained from the Optum Clinformatics Data Mart dataset demonstrate that the methodology is effective and adaptable to large datasets, making it a promising solution for future healthcare data standardization efforts.
@@
Future research will focus on four key areas: (1) extending the framework to accommodate diverse data structures across multiple clinical databases, improving its cross-platform applicability; (2) enhancing the system’s contextual understanding through domain-specific improvements to both the embeddings and re-ranking components; (3) streamlining the validation workflow through improved user interfaces and synthetic training data generation; and (4) establishing a multi-annotator benchmark to reduce bias and quantify agreement (e.g., Cohen’s $\kappa$), thereby strengthening the validity and generalizability of future evaluations.